Calculate the molecular masses of H2, O2, C12, CO2, CH4, C2H6, C2H4, NH3, CH3OH.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The molecular masses are:
H2⇒ 1 × 2 → 2u
O2⇒ 16 × 2 → 32u
Cl2⇒ 35.5 × 2 → 71 u
CO2⇒ 1 × 12 + 2 × 16 = 12 +32 = 44 u
CH4⇒ 1 × 12 + 4 × 1 = 16 u
C2H6 ⇒n2× 12 + 6 × 1 = 30 u
C2H4⇒ (2 × 12) + (4 × 1) = 28 u
NH3⇒ (1 × 14) + (3 × 1) = 17 u
CH3OH ⇒ 12 + (3 × 1) + 16 + 1 = 32 u

Thanks